OpenStudy (anonymous):

the sum of three consecutive odd integers is 165. What are the integers? integer one: integer two: integer three:

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

lets try guessing again

OpenStudy (anonymous):

29+31+33=?

OpenStudy (anonymous):

nah, way too low lets try some higher ones

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

51+53+55=159 still too small

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

how about 53+55+57 ?

OpenStudy (anonymous):

that one looks pretty good

OpenStudy (anonymous):

yeah its right thanks your smart.

OpenStudy (anonymous):

your math teacher wants you to write n + n+2 + n+4 = 165 and solve for n , but that is stupid just guess and check, it has nothing to do with being smart grind it til you find it

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

got more?